# CLEAR: Character Unlearning in Textual and Visual Modalities

### v3

@@ Abstract
Machine Unlearning (MU) is critical for removing private or hazardous information from deep learning models. While MU has advanced significantly in unimodal (text or vision) settings, multimodal unlearning (MMU) remains underexplored due to the lack of open benchmarks for evaluating cross-modal data removal. To address this gap, we introduce CLEAR, the first open-source benchmark designed specifically for MMU. CLEAR contains 200 fictitious individuals and 3,700 images linked with corresponding question-answer pairs, enabling a thorough evaluation across modalities. We conduct a comprehensive analysis of 11 MU methods (e.g., SCRUB, gradient ascent, DPO) across four evaluation sets, demonstrating that jointly unlearning both modalities outperforms single-modality approaches.
@@
Alexey Dontsov1,6, Dmitrii Korzh1,3, Alexey Zhavoronkin2,4, Boris Mikheev3, Denis Bobkov1,6, Aibek Alanov1,6, Oleg Y. Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,6 1AIRI 2MIPT 3Skoltech 4Sber 5MTUCI 6HSE University Correspondence: dontsov@airi.net; tutubalina@airi.net
@@ 1 Introduction
While MU has emerged as a promising alternative to costly retraining—enabling targeted removal of problematic data—existing methods focus almost exclusively on single-modality models. Recent work has studied unlearning in LLMs Yao et al. (2024c, a); Xing et al. (2024); Zhang et al. (2024a) or vision models Li et al. (2024a); Chen and Yang (2023); Tarun et al. (2021), but unlearning in multi-modal language models remains largely unexplored. This leaves a critical gap: multimodal LLMs (MLLMs), which process both visual and textual data, introduce unique challenges for unlearning. For instance, sensitive information may persist across modalities even after removal from one (e.g., a face linked to a name), and unlearning in one modality could degrade performance in another. Despite these risks, no open benchmarks exist to evaluate MU in multimodal settings.
@@
Recently, Chakraborty et al. (2024) pioneers the investigation of unlearning configurations in visual-language models (VLMs) to mitigate cross-modal safety risks, its experimental framework inherits a critical limitation: the datasets used (e.g., PKU-SafeRLHF Ji et al. (2023), JailBreakV-28K Luo et al. (2024)) were designed for safety alignment and truthfulness evaluation, not machine unlearning (MU). This mismatch conflates safety fine-tuning (suppressing harmful outputs) with targeted data removal (erasing specific knowledge traces), potentially overestimating MU efficacy.
@@
  • We comprehensively evaluate 11 recently proposed MU methods on our dataset and show that leading unimodal MU methods struggle in multimodal setups.
@@
We make all the data publicly available111at https://huggingface.co/datasets/fd28/CLEAR.
@@ 2 Related Work
Machine Unlearning. The concept of machine unlearning was initially presented by Cao and Yang (2015). In general, MU methods Cao and Yang (2015); Dwork et al. (2014); Kurmanji et al. (2024); Neel et al. (2021); Sekhari et al. (2021) remove the impact of specific data points from a trained model without requiring full retraining. The goal is to obtain a model that behaves like the forget data was never part of the training set. Recently, textual unlearning in generative language models has attracted attention. Maini et al. (2024) propose a benchmark named TOFU for textual LLM unlearning, consisting of 200 fictitious author profiles defined by attributes like name, birthplace, parent’s names, occupation, and written books, totalling 4,000 question-answer pairs (20 per author). WMDP Li et al. (2024b) includes 3,668 multiple-choice questions to evaluate and benchmark the unlearning of hazardous knowledge in LLMs. To remove knowledge from generative models, Jang et al. (2023) employ gradient ascent on specific target sequences. Eldan and Russinovich (2023) focus on the particular case of unlearning the Harry Potter books from Llama2-7b. Yao et al. (2023) utilize machine unlearning to address harmful responses and eliminate hallucinations. Yao et al. (2024b) examined the unlearning of 2,000 GitHub code files, 500 books, and 500 academic papers from Yi-6B. However, these studies have been restricted to text-only contexts. Our research investigates the multimodal aspects of unlearning.
@@
Multimodal LLMs Liu et al. (2023) usually comprise a modality encoder, a projection layer aligning features to the language space, and a pre-trained language model. While MLLMs have advanced, multimodal unlearning remains under-explored. Cheng and Amiri (2023) introduce MultiDelete, which separates cross-modal embeddings for unlearning but applies only to encoder-decoder models, limiting its use for decoder-only architectures. EFUF Xing et al. (2024) reduces hallucinations in MLLMs by unlearning. It uses CLIP Radford et al. (2021) to detect hallucinations based on MSCOCO-calibrated thresholds, eliminating manual labelling. The method applies three losses: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to preserve fluency. Single Image Unlearning (SIU) Li et al. (2024a) targets visual concept unlearning in VLLMs while preserving textual knowledge and introduces MMUBench. This benchmark spans 20 concepts with 50+ images each, including real-world figures and cartoon characters. However, these benchmarks are not open-sourced. The closest work to ours is Chakraborty et al. (2024), which investigates safety alignment in VLMs by unlearning harmful content. This study was not designed for the exact unlearning setup; therefore, we bridge this gap by conducting the first comprehensive analysis of MU methods in multimodal settings. Unlike safety alignment, our benchmark focuses on forgetting quality (e.g., inability to recall personas) and cross-modal consistency (e.g., erasing both a face and its biography) while maintaining model utility in real-world tasks, highlighting unique challenges in multimodal MU.
@@ 3.1 Preliminaries
Let $f_{\theta}$ denote the base model with parameters $\theta$. It is trained on (train) dataset $D$, and given the unlearning objective, we want to make our model forget a subset of this dataset $D$, called forget set $D_{F}$. The remaining part of the training dataset is called retain set, and we aim to preserve the model’s performance on this data subset $D_{R}:=D\setminus D_{F}$. Additionally, we utilize a holdout set $D_{H}$ such that $D_{H}\cap D=\emptyset$ to establish a reference for the model’s desired behaviour on $D_{F}$ after the unlearning process. In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behaviour on data that was not involved in the training process. Such forgetting procedure is performed by updating the model $f_{\theta}$ with a particular unlearning method, which results in a new unlearned model $f_{\hat{\theta}}$ with parameters $\hat{\theta}$. Evaluation of $f_{\hat{\theta}}$ on the discussed subsets (or particularly on forget set) is called ”inexact” in contrast to the ”exact” evaluation when we directly compare the performance of the unlearned model with a gold model $g_{\omega}$, trained only on the $D_{R}$.
@@
In multi-modal unlearning, compared to unimodal unlearning, each sample contains multiple modalities, e.g., text and image pairs. Let:
@@
where $m\geq 2$ is the number of modalities. We similarly define the forget set $D_{F}$ of multi-modal samples that we aim to unlearn. The model $f_{\theta}$ is now capable of taking multi-modal input (e.g., a textual prompt plus an image) and producing an output $y$. The same forget and retain goals hold.
@@ 3.2 Methods
replaces original labels in the forget set $D_{F}$ with ”I don’t know” responses while minimizing the loss on the retain set $D_{R}$ Maini et al. (2024). The objective, $L_{\text{IDK}}=L(D_{R},\theta)+L(D_{F}^{\text{idk}},\theta)$, ensures that the model retains performance on $D_{R}$ while aligning predictions on $D_{F}$ with uncertainty-based responses.
@@
$L_{\text{LLMU}}=-L(D_{F},\theta)+L(D_{F}^{\text{idk}},\theta)\\ +\sum_{x,y\in D_{R}}\operatorname{KL}(p_{\theta}(y|x)||p_{\hat{\theta}}(y|x)).$
@@
formulates unlearning as a teacher-student setup, where a student model learns from a fixed teacher Kurmanji et al. (2023). The student is optimized to match the teacher on the retain set $D_{R}$ while deviating on the forget set $D_{F}$. The loss function combines KL-divergence for retention ($L_{R}$), enforced divergence for unlearning ($L_{F}$), and task loss ($L_{\text{task}}$), yielding $L_{\text{SCRUB}}=L_{R}-L_{F}+L_{\text{task}}$.
@@
method applies Direct Preference Optimization (DPO) Rafailov et al. (2023) for MU and MMMU. The model is trained to reduce reliance on undesired information through a loss function combining task performance retention ($L_{\text{task}}$) and a DPO-based loss ($L_{\text{DPO}}$), which penalizes deviations from a reference model $\pi_{ref}$ fine-tuned on $D_{F}^{\text{idk}}$ with ”I don’t know” labels as in IDK-tuning.
@@ 4.1 Dataset Generation Process
The generation of synthetic faces for author profiles in our benchmark is motivated by ethical, technical, and practical considerations (see the complete rationale in Appx. A). Firstly, for each of the 200 authors from the TOFU dataset, we extract their name, age, and ethnicity based on the knowledge provided in the original dataset. Also, we generate a pool of 2000 faces using StyleGAN2 Karras et al. (2020) - an established generative model for face synthesis. Each face is scored with a pre-trained image model to determine age, gender, and ethnicity. Then, for each author, we filter a pool of faces with similar characteristics and select the most appropriate one. We found out from textual information that the age distribution of the authors was highly shifted towards the older age group, so we needed to eliminate the age gap between authors’ profiles and corresponding images. To do this, we used the image editing framework proposed in Bobkov et al. (2024) to shift the visual attributes of the faces to make them older. The final distribution of face and author characteristics is shown in Fig.  2. After matching each author with a face, we used the diffusion model Li et al. (2024c) to synthesize images based on the given face and corresponding prompt (Appx.  C).
Figure 3: The overview of our dataset.
We perform a simple reality check to ensure the quality of generated faces. We use the CLIP ViT-L/14 model, usually considered a visual encoder for most of the VLMs, to get embeddings of these three image sets – our faces, CelebA Liu et al. (2015) and WebFace Yi et al. (2014). Then, we calculate the pairwise FID scores on top of the embeddings of these sets, and we get the following results: FID between our faces and CelebA is 74.4, between our faces and WebFace is 69.2, and between CelebA and WebFace is 62.1. This shows that the distance between our faces and the real-world faces is comparable to the distance between two real-world face datasets. Besides the author’s face, the diffusion model needs a textual prompt to produce an image. We ask GPT-4 to generate these prompts from a question-answer pair from TOFU about an author. We generate 8 images for each prompt, evaluate them using an ensemble of fake-detection models, and select the most realistic one. Additionally, GPT-4o generates captions for each (image, visual prompt) pair, which are then included in the dataset to form pairs (image, caption). However, due to restrictions caused by GPT guard breaks and the identification of several bugs in the TOFU dataset (such as a nameless author), the final dataset includes fewer images than text pairs (3,770 compared to 4,000).
@@ 5.1 Evaluation Metrics
We conduct a comprehensive evaluation using ROUGE-L, Probability Score, Truth Ratio, and Forget Quality metrics to thoroughly assess unlearning performance across textual, visual, and multimodal domains. Following Maini et al. (2024); Li et al. (2024a); Xing et al. (2024), this evaluation setup ensures that we capture the effectiveness of unlearning algorithms while examining both retention and forgetfulness of information within models.
@@
Forget Quality calculates the "distance" of the unlearned model to the gold model, which is a proxy metric for the quality of exact unlearning. Following Maini et al. (2024), we take the Truth Ratios distribution of both models. But, instead of the p-value of the Kolmogorov-Smirnov test, we calculate the Jensen-Shannonn distance between these distributions. The latter metric better captures the differences between models, which we additionally check and describe in the Appx. E. To maintain the higher - the better convention, we subtract the distance from 1.

### v4

@@ Abstract
Machine Unlearning (MU) is critical for removing private or hazardous information from deep learning models. While MU has advanced significantly in unimodal (text or vision) settings, multimodal unlearning (MMU) remains underexplored due to the lack of open benchmarks for evaluating cross-modal data removal. To address this gap, we introduce CLEAR, the first open-source benchmark designed specifically for MMU. CLEAR contains 200 fictitious individuals and 3,700 images linked with corresponding question-answer pairs, enabling a thorough evaluation across modalities. We conduct a comprehensive analysis of 11 MU methods (e.g., SCRUB, gradient ascent, DPO) across four evaluation sets, demonstrating that jointly unlearning both modalities outperforms single-modality approaches. The dataset is available at https://huggingface.co/datasets/therem/CLEAR.
@@
Alexey Dontsov2,1, Dmitrii Korzh1,3, Alexey Zhavoronkin4,6, Boris Mikheev3, Denis Bobkov1,2, Aibek Alanov1,2, Oleg Y. Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,4,7 1AIRI 2HSE University 3Skoltech 4Sber AI 5MTUCI 6MIPT 7ISP RAS Research Center for Trusted AI Correspondence: dontsov@airi.net; tutubalina@airi.net
@@ 1 Introduction
MU has emerged as a promising alternative to costly retraining, existing methods focus almost exclusively on single-modality models. Recent work has studied unlearning in LLMs Yao et al. (2024c, a); Xing et al. (2024); Zhang et al. (2024a) or vision models Li et al. (2024a); Chen and Yang (2023); Tarun et al. (2021), but unlearning in multi-modal language models remains largely unexplored. This leaves a critical gap: multimodal LLMs (MLLMs), which process both visual and textual data, introduce unique challenges for unlearning. For instance, sensitive information may persist across modalities even after removal from one (e.g., a face linked to a name), and unlearning in one modality could degrade performance in another. Despite these risks, no open benchmarks exist to evaluate MU in multimodal settings.
@@
Recently, Chakraborty et al. (2024) pioneers the investigation of unlearning configurations in visual-language models (VLMs) to mitigate cross-modal safety risks, its experimental framework inherits a critical limitation: the datasets used in this study (e.g., PKU-SafeRLHF Ji et al. (2023), JailBreakV-28K Luo et al. (2024)) were designed for safety alignment and truthfulness evaluation, not machine unlearning (MU). This mismatch conflates safety fine-tuning (suppressing harmful outputs) with targeted data removal (erasing specific knowledge traces), potentially overestimating MU efficacy.
@@
  • We comprehensively evaluate 11 recently proposed MU methods on our dataset and show that the leading unimodal MU methods struggle in multimodal setups.
@@
We make all data publicly available111 https://huggingface.co/datasets/therem/CLEAR.
@@ 2 Related Work
Machine Unlearning. The concept of machine unlearning was initially presented by Cao and Yang (2015). In general, MU methods Cao and Yang (2015); Dwork et al. (2014); Kurmanji et al. (2024); Neel et al. (2021); Sekhari et al. (2021) remove the impact of specific data points from a trained model without requiring full retraining. The goal is to obtain a model that behaves like the forget data was never part of the training set. Recently, textual unlearning in generative language models has attracted attention. Maini et al. (2024) propose a benchmark named TOFU for textual LLM unlearning, consisting of 200 fictitious author profiles defined by attributes like name, birthplace, parent’s names, occupation, and written books, totaling 4,000 question-answer pairs (20 per author). WMDP Li et al. (2024b) includes 3,668 multiple-choice questions to evaluate and benchmark the unlearning of hazardous knowledge in LLMs. To remove knowledge from generative models, Jang et al. (2023) employ gradient ascent on specific target sequences. Eldan and Russinovich (2023) focus on the particular case of unlearning the Harry Potter books from Llama2-7b. Yao et al. (2023) utilize machine unlearning to address harmful responses and eliminate hallucinations. Yao et al. (2024b) examined the unlearning of 2,000 GitHub code files, 500 books, and 500 academic papers from Yi-6B. However, these studies have been restricted to text-only contexts. Our research investigates the multimodal aspects of unlearning.
Figure 2: The overview of our multimodal dataset. The dataset consists of four sets: retain set, forget set, real faces (knowledge of related concepts such as faces), and real world (to evaluate general visual capabilities).
@@
Multimodal LLMs Liu et al. (2023) usually comprise a modality encoder, a projection layer aligning features to the language space, and a pre-trained language model. While MLLMs have advanced, multimodal unlearning remains under-explored. Cheng and Amiri (2023) introduce MultiDelete, which separates cross-modal embeddings for unlearning but applies only to encoder-decoder models, limiting its use for decoder-only architectures. EFUF Xing et al. (2024) reduces hallucinations in MLLMs by unlearning. It uses CLIP Radford et al. (2021) to detect hallucinations based on MSCOCO-calibrated thresholds, eliminating manual labeling. The method applies three losses: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to preserve fluency. Single Image Unlearning (SIU) Li et al. (2024a) targets visual concept unlearning in VLLMs while preserving textual knowledge and introduces MMUBench. This benchmark spans 20 concepts with 50+ images each, including real-world figures and cartoon characters. However, these benchmarks are not open-sourced. The closest work to ours is Chakraborty et al. (2024), which investigates safety alignment in VLMs by unlearning harmful content. This study was not designed for the exact unlearning setup; therefore, we bridge this gap by conducting the first comprehensive analysis of MU methods in multimodal settings. Unlike safety alignment, our benchmark focuses on forgetting quality (e.g., inability to recall personas) and cross-modal consistency (e.g., erasing both a face and its biography) while maintaining model utility in real-world tasks, highlighting unique challenges in multimodal MU.
@@ 3.1 Preliminaries
Let $f_{\theta}$ denote the base model with parameters $\theta$. It is trained on (train) dataset $D$, and given the unlearning objective, we want to make our model forget a subset of this dataset $D$, called forget set $D_{F}$. The remaining part of the training dataset is called retain set, and we aim to preserve the model’s performance on this data subset $D_{R}:=D\setminus D_{F}$. Additionally, we utilize a holdout set $D_{H}$ such that $D_{H}\cap D=\emptyset$ to establish a reference for the model’s desired behaviour on $D_{F}$ after the unlearning process. In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behavior on data that was not involved in the training process. Such forgetting procedure is performed by updating the model $f_{\theta}$ with a particular unlearning method, which results in a new unlearned model $f_{\hat{\theta}}$ with parameters $\hat{\theta}$. Evaluation of $f_{\hat{\theta}}$ on the discussed subsets (or particularly on forget set) is called ”inexact” in contrast to the ”exact” evaluation when we directly compare the performance of the unlearned model with a gold model $g_{\omega}$, trained only on the $D_{R}$.
@@
In multi-modal unlearning, compared to unimodal unlearning, each sample contains multiple modalities, e.g., text and image pairs, and at least two modalities are processed for the MMU. Let:
@@
where $m\geq 2$ is the number of modalities. We similarly define the forget set $D_{F}$ of multi-modal samples that we aim to unlearn. The model $f_{\theta}$ is now capable of taking multi-modal input (e.g., a textual prompt plus an image) and producing an output $y$. The same forget and retain goals hold. Note that multimodal data allows unimodal unlearning (for example, one can mask tokens of omitted modalities), but multimodal unlearning requires processing of several modalities. From the practical point of view, both MU and MMU depend on optimization loss functions, which at the same time depend on the inputs, either unimodal or multimodal. Some methods, originally proposed for the MU, can naturally be applied for the MMU; however, some MU methods, for example RMU or $\text{SCRUB}_{\text{bio}}$ are non-trivial to be adapted for the MMU.
@@ 3.2 Methods
replaces original labels in the forget set $D_{F}$ with ”I don’t know” responses while minimizing the loss on the retain set $D_{R}$ Maini et al. (2024). The objective $L_{idk}$ ensures that the model retains performance on $D_{R}$ while aligning predictions on $D_{F}$ with uncertainty-based responses:
$L_{idk}=L(D_{R},\theta)+L(D_{F}^{idk},\theta)$
@@
$L_{\text{LLMU}}=-L(D_{F},\theta)+L(D_{F}^{\text{idk}},\theta)\\ +\sum_{x,y\in D_{R}}\operatorname{KL}(p_{\theta}(y|x)||p_{\hat{\theta}}(y|x))$
@@
formulates unlearning as a teacher-student setup, where a student model learns from a fixed teacher Kurmanji et al. (2023). The student is optimized to match the teacher on the retain set $D_{R}$ while deviating on the forget set $D_{F}$. The loss function combines KL-divergence for retention ($L_{R}$), enforced divergence for unlearning ($L_{F}$), and task loss ($L_{\text{task}}$):
$d(x,w^{s})=\operatorname{KL}(p(f(x;w^{o}))||p(f(x;w^{s}))),$
$L_{R}=\frac{\alpha}{|D_{R}|}\sum_{x_{r}\in D_{R}}d(x_{r},w^{s}),$
$L_{F}=\frac{1}{|D_{F}|}\sum_{x_{f}\in D_{F}}d(x_{f},w^{s}),$
$L_{\text{task}}=\frac{\gamma}{|D_{R}|}\sum_{x_{r}\in D_{R}}l(x_{r},y_{r}),$
$L_{\text{SCRUB}}=L_{R}-L_{F}+L_{\text{task}}$
where $f(x;w^{o})$ is the original teacher model with weights $w^{o}$, which are kept unchanged, $f(x;w^{s})$ is the unlearned student model with parameters $w^{s}$, which are optimized.
@@
method applies Direct Preference Optimization (DPO) Rafailov et al. (2023) for MU and MMMU. The model is trained to reduce reliance on undesired information through a loss function combining task performance retention ($L_{\text{task}}$) and a DPO-based loss ($L_{\text{DPO}}$), which penalizes deviations from a reference model $\pi_{ref}$ fine-tuned on $D_{F}^{\text{idk}}$ with “I don’t know” labels as in IDK-tuning.
@@ 4.1 Dataset Generation Process
The generation of synthetic faces for author profiles in our benchmark is motivated by ethical, technical, and practical considerations (see the complete rationale in Appx. A). Firstly, for each of the 200 authors from the TOFU dataset, we extract their name, age, and ethnicity based on the knowledge provided in the original dataset. Also, we generate a pool of 2000 faces using StyleGAN2 Karras et al. (2020) - an established generative model for face synthesis. Each face is scored with a pre-trained image model to determine age, gender, and ethnicity. Then, for each author, we filter a pool of faces with similar characteristics and select the most appropriate one. We found out from textual information that the age distribution of the authors was highly shifted towards the older age group, so we needed to eliminate the age gap between authors’ profiles and corresponding images. To do this, we used the image editing framework proposed in Bobkov et al. (2024) to shift the visual attributes of the faces to make them older. The final distribution of face and author characteristics is shown in Fig.  4. After matching each author with a face, we used the diffusion model Li et al. (2024c) to synthesize images based on the given face and corresponding prompt (Appx.  C).
We perform a simple reality check to ensure the quality of generated faces. We use the CLIP ViT-L/14 model, which is usually considered a visual encoder for most VLMs, to get embeddings of these three image sets – our faces, CelebA Liu et al. (2015) and WebFace Yi et al. (2014). Then, we calculate the pairwise FID scores on top of the embeddings of these sets, and we get the following results: FID between our faces and CelebA is 74.4, between our faces and WebFace is 69.2, and between CelebA and WebFace is 62.1. This shows that the distance between our faces and the real-world faces is comparable to the distance between two real-world face datasets. In addition to the author’s face, the diffusion model needs a textual prompt to produce an image. We ask GPT-4 to generate these prompts from a question-answer pair from TOFU about an author. We generate $8$ images for each prompt, evaluate them using an ensemble of fake detection models, and select the most realistic one. Additionally, GPT-4o generates captions for each (image, visual prompt) pair, which are then included in the dataset to form pairs (image, caption). However, due to restrictions caused by GPT guard breaks and the identification of several bugs in the TOFU dataset (such as a nameless author), the final dataset includes fewer images than text pairs (3,770 compared to 4,000).
@@ 5.1 Evaluation Metrics
We conduct a comprehensive evaluation using ROUGE-L, Probability Score, Truth Ratio, and Forget Quality metrics to thoroughly assess unlearning performance across textual, visual, and multimodal domains. Following Maini et al. (2024); Li et al. (2024a); Xing et al. (2024), this evaluation setup ensures that we capture the effectiveness of unlearning algorithms while examining both the retention and forgetfulness of information within the models.
@@
Forget Quality calculates the “distance” of the unlearned model to the gold model, which is a proxy metric for the quality of exact unlearning. Following Maini et al. (2024), we take the Truth Ratios distribution of both models. However, instead of the p-value of the Kolmogorov-Smirnov test, we calculate the Jensen-Shannonn distance between these distributions. The latter metric better captures the differences between models, which we additionally check and describe in Appx. E. To maintain the higher - the better convention, we subtract the distance from 1.
5.2 Implementation
For the source model, we use LLaVa model Liu et al. (2023) with ViT Dosovitskiy et al. (2021) as visual encoder and LLaMa2-7B Touvron et al. (2023) as a language model. First, we finetune it on the image captioning task using the full CLEAR, both visual and textual parts. We call this model “base”, as it contains forget and retain sets. Then, we perform the unlearning process on it. We use the same hyperparameters for each method. We evaluate the unlearned model based on our metrics from Sec. 5, using the Multi-choice VQA task for the probability score and the image captioning task for the Truth Ratio. For comparison, we also present the metrics of the “gold” model. Experimental results and metrics are shown in Tab. 1, with details provided in Appx. H. In addition, we perform experiments with the QwenVL2 model (Appx. I).
6 Results
  1. RQ1:
  2. RQ2:
  3. RQ3:
6.1 Transferability from Single Domain
To investigate RQ1, we analyze how well the performance of unlearning methods in single modalities predicts their effectiveness in multimodal settings. For the textual domain, we use the TOFU benchmark; for the visual, we use a standard U-MIA approach to the data, consisting of the faces from our dataset; the details for the pipelines and full results are provided in Appx. F and G. For multimodal unlearning, we use our benchmark.